OpenStudy (anonymous):

How do you find an equation of a line tangent to the graph of E when given a point?

ganeshie8 (ganeshie8):

since u already have a `point`, if u could find the `slope` somehow, you can write the tangent line ins `point-slope` form

ganeshie8 (ganeshie8):

taking the derivative of a function at a point gives u the slope at that point : slope = derivative

OpenStudy (anonymous):

Okay, so if I have the point (0,1), how do I find the slope (derivative) at that point?

ganeshie8 (ganeshie8):

where is the graph/function ?

OpenStudy (anonymous):

It's called "E", or the basic graph of an exponential function. It intersects the X-axis at (0,1).

ganeshie8 (ganeshie8):

ahh ok :) so your function is : \(\large y = e^x\)

OpenStudy (anonymous):

exactly

ganeshie8 (ganeshie8):

you dont want to use the derivative formula ? derivative of \(e^x\) is \(e^x\) itself : \(\dfrac{d}{dx}(e^x) = e^x\)

ganeshie8 (ganeshie8):

since you want the slope at (0,1), plugin x = 0 derivative of \(e^x\) at x = 0 is : \(e^0 = 1\)

ganeshie8 (ganeshie8):

so your slope at (0, 1) is `1`

OpenStudy (anonymous):

Huh, well that's easy. I have another problem that says that the X value of a point is 2, could you walk me through that real quick? Wouldn't it be \[e ^{2}=2e ?\]

ganeshie8 (ganeshie8):

\(e^2 = e^2 \ne 2e\)

ganeshie8 (ganeshie8):

Is this the qyestion ? Given the function : \(y = e^x\) find the tangent line at \(x = 2\) ?

OpenStudy (anonymous):

It gives me 2 as the abscissa, not the point which makes it trickier. Otherwise, yes.

ganeshie8 (ganeshie8):

you can get the y-coordinate of point by plugging in x = 2 into ur equation

OpenStudy (anonymous):

Right.

ganeshie8 (ganeshie8):

y = e^x plugin x = 2 : y = e^2 so ur point is (2, e^2)

ganeshie8 (ganeshie8):

since the derivative of \(e^x\) is \(e^x\) itself, slope = derivative = \(e^2\)

ganeshie8 (ganeshie8):

you have hte point and slope, you can write the equation of line

ganeshie8 (ganeshie8):

point : \((2, e^2)\) slope : \(e^2\)

ganeshie8 (ganeshie8):

And yes, \(e\) is just a number.

OpenStudy (anonymous):

Oooh, I got it! Thank you very much for your help my friend. Haha, yeah, my team drilled that all class XD

ganeshie8 (ganeshie8):

u wlc :)